# A semi-supervised framework for diverse multiple hypothesis testing scenarios

### v2

@@ 7 Discussion
$\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in q(B))$
@@
[…]
It follows that:
$\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})$ $=\mathbb{P}(p_{i}\in[0,a)\mid W_{i})$
$\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W_{i})=\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W,\mathbf{x},L_{-i}).$
S.3 RESET controls the FDR or FDP
In this section we prove that given Assumption 3 RESET controls the FDR or the FDP. We use the target-decoy terminology to refer to a positive label $L_{i}=1$ as a target win or simply a target, and a negative label $L_{i}=-1$ as a decoy win or decoy.
Lemma S1.
Let $\tilde{L}_{i}$ for $i\in[p]$ denote the pseudo labels, where $\tilde{L}_{i}=-1$ for a training decoy and $\tilde{L}_{i}=1$ for a pseudo-target. Let $\mathcal{F}$ be the $sigma$-algebra generated by the winning scores $W$, the side information $\mathbf{x}$, and the labels of the false nulls. Let $N$ be the indices of the true nulls, and $L_{-i},\tilde{L}_{-i}$ be the corresponding labels without the $i$th one. Then under Assumption 3, we have for $i\in N$, $L_{i}=\pm 1$ with
$\mathbb{P}(L_{i}=+1\mid\tilde{L}_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})},\quad\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=1)\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})},$
independently of $\mathcal{F},\tilde{L}_{-i},L_{-i}$.
Proof.
By Assumption 3, we know $\{L_{i}\,:\,i\in N\}$ are i.i.d with $P(L_{i}=1)\leq c_{0}$, independently of $\mathcal{G}:=\sigma(\mathcal{F},\tilde{L}_{-i},L_{-i})$. Hence, the following probability statements are independent of $\mathcal{G}$. In addition, since each decoy is randomly assigned to the training decoy set independently of anything else with probability $s$, for each $i\in N$ we have:
$\geq(1-c_{0})\cdot(1-s),$
Hence,
$=\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot\mathbb{P}(L_{i}=-1)}$
$\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}.$
The proof is complete noting that $\frac{c_{0}}{1-s\cdot(1-c_{0})}+\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}=1$ and $L_{i}\in\{\pm 1\}$ a.s.. ∎
Remark 1.
Lemma S1 essentially states that the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are i.i.d. random variables with $\mathbb{P}(L_{i}=+1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ independently of all other information: the scores $W$, the side information $\mathbf{x}$, the labels of the false null discoveries, and all the labels $\tilde{L}$.
See 1
Proof.
We rely on Theorem 3 of Barber and Candès [1] applied to the set of pseudo-targets ($\tilde{L}_{i}=1$) ordered in decreasing order of the learned scores $\tilde{W}_{i}$. Specifically, we assign to each pseudo-target a one-bit p-value:
$\tilde{p}_{i}=\begin{cases}\frac{c_{0}}{1-s\cdot(1-c_{0})}&L_{i}=+1\\ 1&L_{i}=-1\end{cases}.$
Recall that the scores $\tilde{W}$ are themselves a function of the values $(\tilde{L},W,x)$ and clearly the one bit p-values are a function of the labels $L_{i}$. Then it follows from the last Lemma that given the scores $\tilde{W}$ and the one-bit p-values of the false nulls, the one-bit p-values of the true nulls hypotheses ($i\in N$ with $\tilde{L}_{i}=1$) are i.i.d. with $\mathbb{P}(\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})})=\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$, i.e., they stochastically dominate the uniform distribution and hence are valid p-values. Therefore, the condition of Theorem 3 holds so applying, as RESET does, Selective SeqStep+ with $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ to these p-values controls the FDR at level $\alpha$. Note that the original formulation of Selective SeqStep+ is equivalent to Algorithm S1 since we identify the event $\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ with $L_{i}=1$ and $\tilde{p}_{i}=1$ with $L_{i}=-1$. ∎
See 2
Proof.
The proof of Theorem 1 shows that conditional on the learned scores $\tilde{W}$ and the false null labels the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are distributed as i.i.d. ±$\pm 1$ RVs, with $\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$. Assuming for a moment that $\mathbb{P}(L_{i}=1)=\frac{c_{0}}{1-s\cdot(1-c_{0})}$, then RESET satisfies the assumption of the multiple decoy version of FDP-SD with $c=\lambda=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ and therefore it controls the FDP at $\alpha$ with confidence $1-\gamma$ (Section 4 of [40]). In the case that $\mathbb{P}(L_{i}=1)<\frac{c_{0}}{1-s\cdot(1-c_{0})}$, then the procedure is conservative with $c=\lambda=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ and the proof in [40] follows without any modification. ∎
S.4 Further implementation details of RESET
We point out some additional minor details regarding our implementation of RESET. We use the randomxForest package, nnet and mgcv to implement the random forest, two-layer neural network and generalized additive model in R, respectively. We used the default parameters for the random forest, the default parameters for the two-layer neural network and the default parameters for the generalized additive model (except we use drop.intercept = TRUE). If the number of side information variables used by RESET is $\leq 3$, a smoothing spline is fitted with mgcv::s otherwise a natural cubic spline is fitted using 5 degrees of freedom with splines::ns on each side information variable separately (for computational efficiency). For the use of GAM in RESET, we used tryCatch to switch to a random forest in the rare case that the GAM was not able to complete. Some real datasets contain p-values that are ‘zero’. Hence when we calculate $W_{i}=\lvert\Phi^{-1}(p_{i})\rvert$ for such p-values, we obtain Inf in R. We replace instances of Inf for the maximum real-valued score in the dataset. Lastly we used the default settings of RESET for all numerical and real data experiments unless otherwise stated in the main text.
S.5 Further implementation details of Adaptive Knockoffs
@@
There were some small differences between the implementations in the adaptiveKnockoff package in R and the code used in the original manuscript. Accordingly, we used the updated package version. We also fixed a minor bug, which pruned the candidate set $\mathcal{S}$ before checking the estimated FDR is $\leq\alpha$ — possibly missing out on a marginally larger discovery list. There are several parameters that may be set before applying the Adaptive Knockoff filters, e.g., the initial proportion of hypotheses revealed, reveal_prop. We used the same parameter settings in Simulation 1 as Simulation 1 in [44], and the same parameter settings in Simulation 2-4 as their Simulation 2 [44].
@@
In the application to peptide detection, the primary scores $W$ may be negative. Since Adaptive Knockoffs encodes the labels $L$ by using the sign of the scores $W$, we shifted our peptide scores so that the minimum score was zero. For AdaKO EM, any peptide score $\leq 10^{-3}$ was subsequently assigned a zero score. Finally we used, reveal_prop = 10% which reveals the hypotheses that are less than or equal to the 10% of positive scoring hypotheses (which is the same setting used by the simulations). All other parameters were default.
S.6 Further implementation details of the AdaPT methods
@@
We implemented AdaPT using the adaptMT package in R. In Simulation 5 and 6, we used the same settings considered by the authors [33], i.e., we looked at AdaPT GAM in Simulation 5 and AdaPT GLMnet in Simulation 6 with the same parameters. Moreover, we considered AdaPT GLMnet in Simulation 5 with the same settings from Simulation 6, demonstrating the importance of selecting the correct working model. In the real datasets looked at by Zhang et al. [51], we used AdaPT GAM to fit a natural cubic spline using splines::ns with 5 degrees of freedom (knots chosen by default) on each side information variable. The use of natural cubic splines is a suggested option from the package documentation. There was an exception to this for the GTEx datasets, which flagged an error when we tried to apply the spline basis transformation on the ‘chromatin state of the SNP’. In this case, we applied no transformation to the affected side information variable. Zhang et al. mention that they employed a ‘5-degree for each dimension’, but it is not clear to us if they mean the degree of the piece-wise polynomials that make the spline, or the degrees of freedom of the spline. In the latter case, it is not clear how the knots are chosen, or how they overcame the error associated with the ‘chromatin state of the SNP’. Regardless, our results using AdaPT appear to be essentially the same. Moreover, Zhang et al. omits the application of AdaPT to the two fMRI datasets on account of the categorical side information variable used in these datasets. However, Chao and Fithian appear to still apply AdaPTg without problem, and so analogously we applied AdaPT as well. Lastly, in the gene-drug response data, we followed the same implementation by Lei and Fithian [33], which used AdaPT GLM and a collection of candidate side information transformations using splines::ns with degrees of freedom ranging from 6 to 10.
S.7 Further implementation details of the AdaPTg methods
@@
We implemented AdaPTg using the repository https://github.com/patrickrchao/adaptMT which has all the same implementations as AdaPT while offering the ability to define asymmetric regions for mirroring the p-values. Accordingly, we followed the same implementation details as the previous section used by AdaPT and set the so-called ‘masking parameters’ as default.
S.8 Further implementation details of AdaPT-GMMg
@@
We implemented AdaPT-GMMg from the repository https://github.com/patrickrchao/AdaPTGMM. In Simulations 5, we used the GAM filter (model_type = ‘mgcv’) to fit a smoothing spline using mgcv::s on the side information for the M-Step of the EM algorithm. In Simulation 6, we used a generalized linear model filter with regularization (model_type = ‘glmnet’) applied directly to the side information variables. For the real datasets used by Zhang et al. [51], we used the exact same implementation as Chao et al. from the file https://github.com/patrickrchao/AdaPTGMM_Experiments/blob/main/AdaFDR_experiments/run_all_exp.R. In the gene-drug response data, since only one-dimensional side information is used, we copied the parameters selected from the previous real datasets that also considered just one-dimensional side information.
S.9 Further implementation details of ZAP
@@
We used the default parameters of zap_asymp from the ZAP package https://github.com/dmhleung/zap. In all simulations and real data experiments, we considered a natural cubic spline basis expansion on each side information variable with six degrees of freedom using splines::ns. This choice was based on their own analysis on two of the real datasets that we consider here, Airway and Bottomly. The exception was in Simulation 6, where we directly used all 100 side information variables rather than applying a spline basis expansion on each of them. In Simulation 5 and 6, we used the test statistics obtained as input, while in the real data experiments we converted the p-values to z-values by using the transformation ±$\pm\Phi^{-1}(p_{i}/2)$ on each p-value $p_{i}$ where the signs are chosen i.i.d. uniformly.
S.10 Further implementation details of AdaFDR

### v3

@@ C.3 Competition-based multiple testing: the knockoff filter
Once the knockoffs have been constructed, each variable is given a user-defined score $W_{i}([X,\tilde{X}],y)\in\mathbb{R}$ and a label $L_{i}([X,\tilde{X}],y)\in\{\pm 1\}$. As in the competition framework described in Section 2, the idea is that the label $L_{i}$ indicates whether the variable $X_{i}$ or the knockoff $\tilde{X}_{i}$ has greater evidence for being relevant to the model, and the score $W_{i}$ indicates by how much. Commonly, the score and label are combined by setting $sign(W_{i})=L_{i}$; however, given our presentation of RESET, it is instructive to separate them.
In this paper, we consider the Lasso Coefficient Difference (LCD) score, which is the difference in the magnitude of the coefficients of the Lasso regression (Tibshirani (1996)). Specifically, cross-validation is used to determine the $\lambda$ penalty when regressing on the augmented design matrix $X,\tilde{X}$]. We denote the absolute values of the Lasso($\lambda$)-fitted coefficients as $Z_{i}$ for the variables and $\tilde{Z}_{i}$ for the knockoffs. The scores and labels are then defined as
@@
where hypotheses with zero scores are thrown out.
C.4 Further implementation details of Adaptive Knockoffs
@@
There were some small differences between the implementations in the adaptiveKnockoff package in R and the code used in the original manuscript. Accordingly, we used the updated package version. We also fixed a minor bug, which pruned the candidate set $\mathcal{\hat{S}}_{t}$ before checking the estimated FDR is $\leq\alpha$ — possibly missing out on a marginally larger discovery list. There are several parameters that may be set before applying the Adaptive Knockoff filters, e.g., the initial proportion of hypotheses revealed, reveal_prop. We used the same parameter settings in Simulation 1 here as in Simulation 1 in Ren and Candès (2023) (which uses one-dimensional side information), and the same parameter settings in Simulations 2-4 here as Simulation 2 in Ren and Candès (2023) (which uses two-dimensional side information or more).
@@
In the application to peptide detection, the primary score $W$ can be negative. Since Adaptive Knockoffs encodes the labels $L$ by using the sign of the scores $W$, we shifted our peptide scores so that the minimum score was zero in those cases. For AdaKO EM, any peptide score $\leq 10^{-3}$ was subsequently assigned a zero score. Finally we used, reveal_prop = 10% which reveals the hypotheses that are less than or equal to the 10% of positive scoring hypotheses (which is the same setting used by the simulations). All other parameters were default.
C.5 Simulation details for the competition setting
C.5.1 Simulation 1: Linear model with one-dimensional side information
Our first simulation is based on Simulation 1 of Ren and Candès (2023). Specifically, $n=1000$ independent observations from the joint data distribution of $(X,Y)$ are generated by drawing $X$ from a hidden markov model (HMM) with $p=900$ variables and $Y|X\sim\mathcal{N}(X\beta,I)$ (linear model). The $\beta$’s are selected in the following way. Let $k$ denote the number of nonzero $\beta$’s. Then $k$ indices from $\{1,2,\dots,2k\}$ are sampled without replacement, each with probability proportional to $1/i^{2}$. The $k$ chosen indices are the nonzero $\beta$’s and are assigned a value of $\pm 3.5/\sqrt{n}$, where the signs are chosen i.i.d. uniformly. We consider $k\in\{50,150,300\}$ while the original simulation only considers $k=150$. We generate HMM knockoffs from Sesia et al. (2019). Each hypothesis is supported by the side information $\mathbf{x}_{i}=i$, which should help in detecting the nonzero $\beta$’s, because smaller values of $i$ are more likely to correspond to a nonzero $\beta_{i}$.
We generated 100 independent datasets each consisting of drawing $n$ observations as described above. The original simulation in Ren and Candès (2023) sampled the $\beta$’s once and fixed them across the 100 datasets. Here we redrew the $\beta$’s for each dataset so that we may average our results over a range of models, as we noticed that all the methods exhibited significant variability between the draws.
C.5.2 Simulation 2: Linear model with two-dimensional side information
In a second simulation, we considered a larger linear model where $n=2000$ observations are taken from the joint distribution $(X,Y)$ as described in Section C.5.1, with $p=1800$. Accordingly, we also adjusted the number of false nulls to $k\in\{150,300,450\}$. Note that unlike the previous simulation, it is not relevant how the $k$ variable indices are selected from $\{1,2,\dots,p\}$. Each hypothesis is equipped with the following two-dimensional side information. For the $i$th hypothesis, we draw a pair of observations $\mathbf{x}_{i}$ from a bivariate normal $\mathcal{N}(\boldsymbol{\mu},I)$ with $\boldsymbol{\mu}=(0,0)$ for a true null hypothesis and $\boldsymbol{\mu}=\pm(2,2)$ for a false null hypothesis, where the signs are chosen i.i.d. uniformly. We generated 20 independent datasets of this larger simulation instead of 100, where each involved drawing those $n$ observations, $\beta$ coefficients, and side information variables.
C.5.3 Simulation 3: Logistic model with two-dimensional side information
Next, we consider Simulation 2 by Ren and Candès (2023). They generate $n=1000$ observations with $p=1600$ variables sampled from the joint data distribution of $(X,Y)$, where $X$ is distributed according to a multivariate normal distribution and $Y|X\sim\text{Bernoulli}\left(e^{\beta\cdot X}/\left(1+e^{\beta\cdot X}\right)\right)$ is a logistic model (see Ren and Candès (2023) for further details). Each hypothesis’ side information is a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from the lattice $20,19]\times[-20,19]\subseteq\mathbb{Z}^{2}$. The nonzero $\beta$’s are chosen according to the position of $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the lattice. An image of the nonzero $\beta$’s is given in Figure C9. Each nonzero $\beta$ is set to $\pm 25/\sqrt{n}$, where the signs are chosen i.i.d. uniformly. We generate approximate SDP Gaussian knockoffs as described in Candès et al. (2018).
We generated 100 independent datasets drawing $n$ observations as above in each but for 2 of those AdaKO EM failed with an error so we used only 98 of the datasets. The original simulation in Ren and Candès (2023) sampled the sign of the $\beta$’s once, fixing them across the 100 datasets. Instead, we redraw the sign of the $\beta$’s in each of the 98 datasets since we found that the results of all the methods significantly varied with those draws.
Figure C9: Location of nonzero $\beta$’s according to the pair of values $(x,y)$ in the lattice. The same image is given in Ren and Candès (2023).
C.5.4 Simulation 4: Large $p$ and $n$ and 3-d side information
Finally, in the last simulation, $n=10K$ observations with $p=10K$ variables and $k=0.15p$ relevant variables are drawn from the joint data distribution $(X,Y)$ exactly as described in Simulation 1. Each variable is equipped with the following three-dimensional side information. The first $k$ variables are randomly assigned a value drawn from $i\in\{1,\dots,2k\}$ taken without replacement and having probability proportional to $1/i^{2}$. The remaining pkp-k variables are uniquely assigned to one of the remaining values from $i\in\{1,2,\dots,p\}$ uniformly at random. Overall, the final assigment is a permutation of the integers in $\{1,\dots,p\}$, subject to the constraint that the first $k$ variables take values exclusively from $\{1,\dots,2k\}$. We expect most of the nonzero $\beta$’s to have small values of $i$, but the effect of the random assignment makes the resulting side information variable less informative than in Simulation 1, as depicted in Figure C10. This process is repeated three times, thus associating three side information variables with each hypothesis, which taken together, should be reasonably informative. We generated 5 independent samples, drawing the $n$ observations, $\beta$ coefficients, and side information in each.
Figure C10: Scatterplot of each hypothesis according to the LCD score denoted as $W$ (x-axis) and side information variable (y-axis). Each hypothesis is coloured according to whether it is a true (red) or false null hypothesis (blue). In (A), the side information variable is the index of the hypothesis in Simulation 1 (Section C.5.1). In (B), we use one of the three side information variables as described in Simulation 4 (Section C.5.4).
C.6 Competition-based multiple testing: peptide detection
Competition-based multiple testing was first established in the proteomics community where the goal is to infer which proteins, i.e., long chains of amino acids, are present in a biological sample (Elias and Gygi (2007)). Direct protein detection is challenging, and so a bottom-up approach is taken where the proteins are digested into small chains of amino acids, called peptides, and the objective is to identify these peptides instead. Then, a process called liquid chromatography tandem mass spectrometry (LC-MS/MS) is used in which a first round of mass spectrometry isolates each sample peptide and reports the abundance or intensity of each sample peptide. In a second round of mass spectrometry, each sample peptide that reports a high enough intensity is selected and further fragmented. The intensity of each fragment is then reported in a corresponding fragmentation spectrum. Each such spectrum can be thought of as a fingerprint of the peptide that generated the spectrum, allowing us to hypothesize which peptides are present in the sample.
The search phase described above is dictated by a range of user-specified parameters. The most common configuration of these parameters correspond to two broad types of searches, namely narrow and open searches, which we consider in this paper. Briefly a ‘narrow’ search, which is the default search mode, only searches each spectrum in the spectrum file against peptides in the target-decoy database that have a theoretical mass within a small tolerance of the mass of the sample peptide that generated the experimental spectrum. On the other hand, an ‘open’ search makes this mass-tolerance larger so that a spectrum may match with a peptide with a completely different mass. This is desirable at times, since sample peptides can undergo post-translational modifications which modify the mass of the peptide so it no longer matches its theoretical mass. However, when searching in an ‘open’ mode, the modified spectrum may still be correctly matched.
The same database peptide, target or decoy, may be matched multiple times with different scores, or have no match at all. Hence, we associate with $H_{i}$ the score $Z_{i}$, defined as the maximum scoring PSM associated with the $i$th target peptide, and by its design the higher the score, the more likely $H_{i}$ is false, i.e., the peptide is in the sample. Similarly, we define the null drawn $\tilde{Z}_{i}$ as the maximum scoring PSM associated to the $i$th target’s corresponding decoy peptide. Target or decoy peptides that have no matching spectra are assigned a score of -$\infty$. Finally, we obtain a winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and a label $L_{i}$ indicating whether the winning score was from the $i$th target or its decoy (where ties are randomly broken). Peptides with a winning score of -$\infty$, indicating that neither the target nor its paired decoy have a matching spectra, are thrown out. Finally, a competition-based multiple testing procedure like SSS+ or FDP-SD may be applied to these scores and labels (Nesvizhskii (2010)).
C.7 Further implementation details of the AdaPT methods
@@
We applied AdaPT using the adaptMT package in R. In Simulations 5 and 6, we used the same settings considered by Lei and Fithian (2018), i.e., we used AdaPT GAM in Simulation 5 and AdaPT GLMnet in Simulation 6 with the same parameters. Moreover, we considered AdaPT GLMnet in Simulation 5 with the same settings from Simulation 6, demonstrating the importance of selecting the correct working model. In Simulations 7 and 8, we used the same settings as Simulation 5. In the real datasets looked at by Zhang et al. (2019), we used AdaPT GAM to fit a natural cubic spline using splines::ns with 5 degrees of freedom (knots chosen by default) on each side information variable. The use of natural cubic splines is a suggested option from the package documentation. There was an exception to this for the GTEx datasets, which flagged an error when we tried to apply the spline basis transformation on the ‘chromatin state of the SNP’. In this case, we applied no transformation to the affected side information variable. Zhang et al. mention that they employed a “5-degree for each dimension”, but it is not clear to us if they mean the degree of the piece-wise polynomials that make the spline, or the degrees of freedom of the spline. In the latter case, it is not clear how the knots are chosen, or how they overcame the error associated with the ‘chromatin state of the SNP’. Regardless, our results using AdaPT appear to be essentially the same. Moreover, Zhang et al. omit the application of AdaPT to the two fMRI datasets on account of the categorical side information variable used in these datasets. However, Chao and Fithian appear to still apply AdaPTg without problem, and so analogously we applied AdaPT as well. Lastly, in the gene-drug response data, we followed the same implementation by Lei and Fithian (2018), which used AdaPT GLM and a collection of candidate side information transformations using splines::ns with degrees of freedom ranging from 6 to 10.
C.8 Further implementation details of the AdaPTg methods
@@
We downloaded AdaPTg from the repository https://github.com/patrickrchao/adaptMT which has all the same implementations as AdaPT while offering the ability to define asymmetric regions for mirroring the p-values. Accordingly, we followed the same implementation details as the previous section used by AdaPT and set the so-called ‘masking parameters’ as default.
C.9 Further implementation details of AdaPT-GMMg
@@
We downloaded AdaPT-GMMg from the repository https://github.com/patrickrchao/AdaPTGMM. In Simulations 5, 7 and 8, we used the GAM filter (model_type = ‘mgcv’) to fit a smoothing spline using mgcv::s on the side information for the M-Step of the EM algorithm. In Simulation 6, we used a generalized linear model filter with regularization (model_type = ‘glmnet’) applied directly to the side information variables. For the real datasets used by Zhang et al. (2019), we used the exact same implementation as Chao et al. from the file https://github.com/patrickrchao/AdaPTGMM_Experiments/blob/main/AdaFDR_experiments/run_all_exp.R. In the gene-drug response data, since only one-dimensional side information is used, we copied the parameters selected from the previous real datasets that also considered just one-dimensional side information.
C.10 Further implementation details of ZAP
@@
We used the default parameters of zap_asymp from the ZAP package https://github.com/dmhleung/zap. In all simulations and real data experiments, we considered a natural cubic spline basis expansion on each side information variable with six degrees of freedom using splines::ns. This choice was based on their own analysis on two of the real datasets that we consider here, Airway and Bottomly. The exception was in Simulation 6, where we directly used all 100 side information variables rather than applying a spline basis expansion on each of them. ZAP requires explicit test-statistics, specifically z-scores, as input. In Simulations 5, 7 and 8, we used the test statistics as input. Since the p-values are obtained from a one-sided test, there is a one-to-one correspondence between the p-value information and test statistic information. Thus our comparisons in Simulations 5, 7, and 8, are still effectively based on the p-values as we addressed in Section 3.1. In Simulation 6, explicit z-score test statistics were not generated by the simulation. Hence, we used the inverse normal CDF $\Phi^{-1}$ to convert the p-values to a z-score test statistic. In the real data experiments, instead of directly using the $\Phi^{-1}(p_{i})$, we converted the p-values to z-values using the transformation $\pm\Phi^{-1}(p_{i}/2)$ on each p-value $p_{i}$ where the signs are chosen i.i.d. uniformly. Otherwise, applying the transformation $\Phi^{-1}(p_{i})$ to true null p-values with a non-decreasing density, or a spike near 1, will produce consistently large z-scores and will be incorrectly discovered by ZAP as an artefact of the data transformation.
C.11 Further implementation details of AdaFDR